# Online Detecting LLM-Generated Texts via Sequential Hypothesis Testing by Betting

### v2

@@ Appendix F Proof of Proposition 2
which indicats that reject the null hypothesis when either $W_{t}^{A}\geq 2/\alpha$ or $W_{t}^{B}\geq 2/\alpha$ is a level-$\alpha$ sequential test. The detection process for the composite hypothesis is shown as Algorithm 3.
We consider a constant $u^{\prime}$ as below, since the score discrepancy $g_{t}\in[-d_{t},d_{t}]$ now becomes $(g_{t}-\epsilon)\in[-d_{t}-\epsilon,d_{t}-\epsilon]$, then
$u^{\prime}=\frac{-\sum_{t=1}^{T}(g_{t}-\epsilon)}{2d^{\prime}_{*}\cdot\left(\sum_{t=1}^{T}(g_{t}-\epsilon)^{2}+\left\lvert\sum_{t=1}^{T}(g_{t}-\epsilon)\right\rvert\right)}\in[-\frac{1}{2d^{\prime}_{*}},\frac{1}{2d^{\prime}_{*}}].$
We denote $d^{\prime}_{*}=d_{*}+\epsilon$, where $d_{*}:=\max\limits_{t}\lvert d_{t}\rvert$. Since $g_{t}\in[-d_{t},d_{t}]$, $u^{\prime}\in[-1/2d^{\prime}_{*},1/2d^{\prime}_{*}]$. The tangent bound needs to be applied $\log(1+a)\geq a-a^{2}$ for $a\in[-1/2,1/2]$ to $-(g_{t}-\epsilon)u^{\prime}$ to get the lower bound of wealth. We have $-(g_{t}-\epsilon)u^{\prime}\in[-(d_{t}-\epsilon)/2d^{\prime}_{*},(d_{t}+\epsilon)/2d^{\prime}_{*}]$, since $0<\epsilon\ll 1$ and $d^{\prime}_{*}\geq d_{t}+\epsilon\geq d_{t}-\epsilon$ for all $t$, then $-(g_{t}-\epsilon)u^{\prime}\in[-1/2,1/2]$ can still hold.
Define
$g_{i}:=\phi(x_{i})-\phi(y_{i}),\quad S^{\prime}_{t}:=\sum_{i=1}^{t}(g_{i}-\epsilon),\quad Q^{\prime}_{t}:=\sum_{i=1}^{t}(g_{i}-\epsilon)^{2}.$
We follow the similar process as before and get
$W_{t}^{A}\gtrsim\frac{1}{{Q^{\prime}}_{t}^{\frac{1}{2\gamma}}}\exp\left(\frac{2{d^{\prime}}_{*}-1}{4{d^{\prime}}_{*}^{2}}\cdot\frac{{S^{\prime}}_{t}^{2}}{{Q^{\prime}}_{t}+\lvert{S^{\prime}}_{t}\rvert}\right)\geq\frac{1}{(t{d^{\prime}}_{*}^{2})^{\frac{1}{2\gamma}}}\exp\left(\frac{2{d^{\prime}}_{*}-1}{4{d^{\prime}}_{*}^{2}}\cdot\frac{{S^{\prime}}_{t}^{2}}{t{d^{\prime}}_{*}^{2}+t{d^{\prime}}_{*}}\right),\quad\forall t\geq 1.$ (63)
It can still give the guarantee of asymptotic power one. As for the expected stopping time, when $P\in H_{1}^{A}$. For the stopping time $\tau>0$, we have
@@
$\mathbb{E}[\tau]=\sum_{t=1}^{\infty}P(\tau>t)\leq\sum_{t=1}^{\infty}P\left(\log(W_{t}^{A})<\log(2/\alpha)\text{ or }\log(W_{t}^{B})<\log(2/\alpha)\right)=2\sum_{t=1}^{\infty}P({E^{\prime}}_{t}),$
where ${E^{\prime}}_{t}=\left\{\log(W_{t}^{A})<\log(2/\alpha)\right\}$.
Based on the first inequality of (63), we have
${E^{\prime}}_{t}$$\subseteq\left\{\log\left(\frac{1}{{Q^{\prime}}_{t}^{\frac{1}{2\gamma}}}\exp\left(\frac{2d_{*}-1}{4{d^{\prime}}_{*}^{2}}\cdot\frac{{S^{\prime}}_{t}^{2}}{({Q^{\prime}}_{t}+\lvert S_{t}\rvert)}\right)\right)<\log(2/\alpha)\right\}$
$\Rightarrow{E^{\prime}}_{t}$$\subseteq\left\{{{S^{\prime}}_{t}^{2}}<\frac{4{d^{\prime}}_{*}^{2}}{2{d^{\prime}}_{*}-1}\left({Q^{\prime}}_{t}+\lvert{S^{\prime}}_{t}\rvert\right)\left(\log(2/\alpha)-\log(1/{Q^{\prime}}_{t}^{\frac{1}{2\gamma}})\right)\right\}$
@@
$\subseteq\left\{{S_{t}^{2}}<\frac{4{d^{\prime}}_{*}^{2}}{2{d^{\prime}}_{*}-1}\left({Q^{\prime}}_{t}+\sum_{i=1}^{t}\lvert g_{i}-\epsilon\rvert\right)\left(\log(2/\alpha)-\log(1/Q_{t}^{\frac{1}{2\gamma}})\right)\right\}.$ (64)
We define ${V^{\prime}}_{t}:=\sum_{i=1}^{t}\lvert g_{i}-\epsilon\rvert$. After applying the Chernoff bound (see e.g., Harvey [2023]) over random variables ${V^{\prime}}_{t}/{d^{\prime}}_{*}\in[0,1]$ and ${Q^{\prime}}_{t}/{d^{\prime}}_{*}^{2}\in[0,1]$, the upper bound on ${V^{\prime}}_{t}$ and ${Q^{\prime}}_{t}$ can be given, which holds for all $t\geq 17$. With a probability of at least $(1-1/t^{2})$, we have ${{V^{\prime}}_{t}}/{{d^{\prime}}_{*}}\leq 2t$. With a probability of at least $(1-1/t^{2})$, ${{Q^{\prime}}_{t}}/{{d^{\prime}}_{*}}\leq 2t$.
Now, ${H^{\prime}}_{t}=\left\{\frac{{Q^{\prime}}_{t}}{{d^{\prime}}_{*}^{2}}\leq 2t\right\}\cap\left\{\frac{{V^{\prime}}_{t}}{{d^{\prime}}_{*}}\leq 2t\right\}$, we get
@@
${E^{\prime}}_{t}\cap{H^{\prime}}_{t}$$\subseteq\left\{{{S^{\prime}}_{t}^{2}}<\frac{4{d^{\prime}}_{*}^{2}}{2{d^{\prime}}_{*}-1}\left(2t{d^{\prime}}_{*}^{2}+2t{d^{\prime}}_{*}\right)\left(\log(2/\alpha)+\log(2t{d^{\prime}}_{*}^{2})^{\frac{1}{2\gamma}}\right)\right\}$
$\subseteq\left\{\left\lvert\frac{{S^{\prime}}_{t}}{{d^{\prime}}_{*}}\right\rvert<\underbrace{\sqrt{\frac{8{d^{\prime}}_{*}({d^{\prime}}_{*}+1)t}{2{d^{\prime}}_{*}-1}\cdot\log\left(2\cdot(2t{d^{\prime}}_{*}^{2})^{\frac{1}{2\gamma}}/\alpha\right)}}_{:={R^{\prime}}}\right\}.$ (65)
All steps are the same as before except for the superscripts, then we apply a Hoeffding’s bound over the independent random variables ${S^{\prime}}_{t}/{d^{\prime}}_{*}\in[-1,1]$ to get
$\left\lvert\frac{{S^{\prime}}_{t}}{{d^{\prime}}_{*}}\right\rvert\geq\left\lvert\mathbb{E}\left[\frac{{S^{\prime}}_{t}}{{d^{\prime}}_{*}}\right]\right\rvert-\sqrt{2t\cdot\log(2t^{2})}\geq\frac{t(\Delta-\epsilon)}{{d^{\prime}}_{*}}-\sqrt{2t\cdot\log(2t^{2})}\geq\frac{t(\Delta-\epsilon)}{{d^{\prime}}_{*}}-\sqrt{4t\cdot\log(2t)},$
where $\Delta=\lvert\mu_{x}-\mu_{y}\rvert$, the second inequality is derived based on the triangle inequality that $\lvert a+b\rvert\leq\lvert a\rvert+\lvert b\rvert$. Then we can rearrange $\lvert(a-b)+b\rvert\leq\lvert a-b\rvert+\lvert b\rvert$ to get $\lvert a-b\rvert\geq\lvert a\rvert-\lvert b\rvert$. Thus,
$\left\lvert\mathbb{E}\left[\frac{{S^{\prime}}_{t}}{{d^{\prime}}_{*}}\right]\right\rvert$ $=\frac{\left\lvert\mathbb{E}\left[\sum_{i=1}^{t}(g_{i}-\epsilon)\right]\right\rvert}{{d^{\prime}}_{*}}$
$=\frac{\left\lvert\mathbb{E}\left[\sum_{i=1}^{t}(\phi({x}_{t})-\phi({y}_{t})-\epsilon)\right]\right\rvert}{d^{\prime}_{*}}$
$=\frac{\left\lvert\sum_{i=1}^{t}\mathbb{E}\left[\phi({x}_{t})-\phi({y}_{t})-\epsilon\right]\right\rvert}{d^{\prime}_{*}}$
$=\frac{\left\lvert t(\mu_{x}-\mu_{y}-\epsilon)\right\rvert}{d^{\prime}_{*}}$
$\geq\frac{t\left\lvert\mu_{x}-\mu_{y}\right\rvert-\epsilon}{d^{\prime}_{*}}\quad(\text{since }\epsilon\text{ is positive}).$
We thus have ${t(\Delta-\epsilon)}/{{d^{\prime}}_{*}}-\sqrt{4t\cdot\log(2t)}\geq R^{\prime}$, or alternatively,
@@
$\frac{t(\Delta-\epsilon)}{{d^{\prime}}_{*}}-\sqrt{4t\cdot\log(2t)}$ $\geq\sqrt{\frac{8{d^{\prime}}_{*}({d^{\prime}}_{*}+1)t}{2{d^{\prime}}_{*}-1}\cdot\log\left(\frac{2\cdot(2t{d^{\prime}}_{*}^{2})^{\frac{1}{2\gamma}}}{\alpha}\right)}.$
Now the remaining steps essentially follow those in Proposition 1. Hence, we can obtain that the expected stopping time satisfies:
@@
$\mathbb{E}[\tau]\lesssim\frac{{d^{\prime}}_{*}^{3}}{(\Delta-\epsilon)^{2}}\cdot\log\left(\frac{{d^{\prime}}_{*}^{(3+\frac{1}{\gamma})}}{(\Delta-\epsilon)^{2}\alpha}\right).$ (66)
We recall that ${d^{\prime}}_{*}=d_{*}+\epsilon$, this completes the proof.
@@
Appendix G Experiment Results of Detecting 2024 Olympic News or Machine-generated News

### v3

@@ Appendix F Proof of Proposition 2
which indicates that rejecting the null hypothesis when either $W_{t}^{A}\geq 2/\alpha$ or $W_{t}^{B}\geq 2/\alpha$ is a level-$\alpha$ sequential test. The detection process for the composite hypothesis is shown as Algorithm 3.
2. Asymptotic Power One. Now let us switch to analyze the test power of the proposed algorithm. We first recall the relation between regret and wealth in the testing-by-betting game (c.f. equation 18), i.e.,
$\ln(W_{t}^{A})=\ln(W_{t}^{A}(\theta_{*}))-\text{Regret}_{t}(\theta_{*}),$ (32)
where $W_{t}^{A}$ is the learner’s wealth at round $t$, $\theta_{*}$ is any benchmark which lies in the decision space $\mathcal{K}_{*}^{\prime}=[-\frac{1}{2d_{*}^{\prime}},0]$ with $d_{*}^{\prime}=d_{*}+\epsilon$.
Denote $\omega_{*}^{A}:=\mathbb{E}[\ln(1-(g-\epsilon)\theta_{*})]$ the expected wealth of the benchmark in a single round. For the notation simplicity, in the following, we will drop the superscript $A$ in $\omega_{*}^{A}$ when it is clear to the context. Taking the expectation on both sides of equation 32, we have
$\mathbb{E}[\ln(W_{t}^{A})]$ $=\mathbb{E}[\ln(W_{t}^{A}(\theta_{*}))]-\mathbb{E}\left[\text{Regret}_{t}(\theta_{*})\right]$
$=\mathbb{E}\left[\sum_{s=1}^{t}\ln(1-(g_{s}-\epsilon)\theta_{*})\right]-\mathbb{E}[\text{Regret}_{t}(\theta_{*})]$
$=t\omega_{*}-\mathbb{E}\left[\text{Regret}_{t}(\theta_{*})\right],$
where the last equality holds by assuming that the random variables $(g_{s})_{s\geq 1}$ are i.i.d.
We now analyze the probability that the null has not been rejected by time $t$, which is when both the events $\{W_{t}^{A}<\frac{2}{\alpha}\}$ and $\{W_{t}^{B}<\frac{2}{\alpha}\}$ hold. We have
$\mathbb{P}\left[W_{t}^{A}<\frac{2}{\alpha}\right]$ $=\mathbb{P}\left[\ln(W_{t}^{A})<\ln\left(\frac{2}{\alpha}\right)\right]$
$=\mathbb{P}\left[\ln(W_{t}^{A})-\mathbb{E}[\ln(W_{t}^{A}(\theta_{*}))]<\ln\left(\frac{2}{\alpha}\right)-\mathbb{E}[\ln(W_{t}^{A}(\theta_{*}))]\right]$
@@
$=\mathbb{P}\left[\ln(W_{t}^{A}(\theta_{*}))-\text{Regret}_{t}(\theta_{*})-\mathbb{E}[\ln(W_{t}^{A}(\theta_{*}))]<\ln\left(\frac{2}{\alpha}\right)-\mathbb{E}[\ln(W_{t}^{A}(\theta_{*}))]\right]$
$=\mathbb{P}\left[\ln(W_{t}^{A}(\theta_{*}))-\mathbb{E}[\ln(W_{t}^{A}(\theta_{*}))]<\text{Regret}_{t}(\theta_{*})+\ln\left(\frac{2}{\alpha}\right)-t\omega_{*}\right].$ (33)
Now we are going to show that $\text{Regret}_{t}(\theta_{*})+\ln\left(\frac{2}{\alpha}\right)-t\omega_{*}\leq\frac{-t\omega_{*}}{2}$ when $t$ is sufficiently large. Then, from equation 33, we will have $\mathbb{P}\left[\ln(W_{t}^{A}(\theta_{*}))-\mathbb{E}[\ln(W_{t}^{A}(\theta_{*}))]<\text{Regret}_{t}(\theta_{*})+\ln\left(\frac{2}{\alpha}\right)-t\omega_{*}\right]\leq\mathbb{P}\left[\ln(W_{t}^{A}(\theta_{*}))-\mathbb{E}[\ln(W_{t}^{A}(\theta_{*}))]<-\frac{t}{2}\omega_{*}\right]$.
From the regret bound of ONS equation 15, it suffices to show that
@@
$\frac{1}{2\gamma}\ln\left(1+\sum_{s=1}^{t}\frac{(g_{s}-\epsilon)^{2}}{(1-(g_{s}-\epsilon)\theta_{s})^{2}}\right)+\frac{\gamma}{2}D_{1}^{2}+\ln\left(\frac{2}{\alpha}\right)\leq\frac{t}{2}\omega_{*}.$ (34)
We will require the time $t$ to satisfy
$t\geq\frac{2\gamma D_{1}^{2}}{\omega_{*}},$ (35)
which leads to $\frac{\gamma}{2}D_{1}^{2}\leq\frac{t\omega_{*}}{4}$. Therefore, to guarantee equation 34, it suffices to find $t$ such that
@@
$\frac{t\omega_{*}}{4}\geq\ln\left(\frac{2}{\alpha}\right)+\frac{1}{2\gamma}\ln\left(1+\sum_{s=1}^{t}\frac{(g_{s}-\epsilon)^{2}}{(1-(g_{s}-\epsilon)\theta_{s})^{2}}\right).$
We further note that $\frac{(g_{s}-\epsilon)^{2}}{(1-(g_{s}-\epsilon)\theta_{s})^{2}}\leq\frac{(d^{\prime}_{*})^{2}}{(1-(d_{s}+\epsilon)/2d_{s})^{2}}\leq c(d^{\prime}_{*})^{2}$, for some constant $c\geq\left(\frac{2d_{s}}{d_{s}-\epsilon}\right)^{2}$, which in turns implies a sufficient condition:
$\frac{t\omega_{*}}{4}\geq\ln\left(\frac{2}{\alpha}\right)+\frac{1}{2\gamma}\ln\left(1+ct(d^{\prime}_{*})^{2}\right).$
The above condition together with equation 35 suggests that $\text{Regret}_{t}(\theta_{*})+\ln\left(\frac{2}{\alpha}\right)-t\omega_{*}\leq\frac{-t\omega_{*}}{2}$ holds when
@@
$t\gtrsim\frac{1}{\omega_{*}}\ln\left(\frac{(d^{\prime}_{*})^{1/\gamma}}{\omega_{*}\alpha}\right).$ (36)
Under equation 36, we hence have
$\mathbb{P}\left[W_{t}^{A}<\frac{2}{\alpha}\right]$ $\leq\mathbb{P}\left[\ln(W_{t}^{A}(\theta_{*}))-\mathbb{E}[\ln(W_{t}^{A}(\theta_{*}))]<-\frac{t}{2}\omega_{*}\right].$ (37)
Denote $\psi_{t}:=\ln(1-(g_{t}-\epsilon)\theta_{*})-\mathbb{E}[\ln(1-(g_{t}-\epsilon)\theta_{*})]\in[\ln(1/2)-\omega_{*},\ln(3/2)-\omega_{*}]$. We have that $\ln(W_{t}^{A}(\theta_{*}))-\mathbb{E}[\ln(W_{t}^{A}(\theta_{*}))]=\sum_{s=1}^{t}\psi_{s}$, which is a sum of zero-mean i.i.d bounded random variables. Specifically, $\psi_{t}$ is a sub-Gaussian with parameter $\sigma:=\frac{1}{2}\ln(3)$. By Hoeffding’s inequality, we have
$\mathbb{P}\left[\frac{1}{t}\sum_{s=1}^{t}\psi_{s}\leq-c\right]\leq\exp\left(-\frac{tc^{2}}{2\sigma^{2}}\right),$ (38)
for any constant $c>0$. Then,
$\mathbb{P}\left[\ln(W_{t}^{A}(\theta_{*}))-\mathbb{E}[\ln(W_{t}^{A}(\theta_{*}))]<-\frac{t}{2}\omega_{*}\right]=\mathbb{P}\left[\frac{1}{t}\sum_{s=1}^{t}\psi_{s}<-\frac{1}{2}\omega_{*}\right]\leq\exp\left(-\frac{(\omega_{*})^{2}}{2(\ln 3)^{2}}t\right),$
where we let $c\leftarrow\frac{\omega_{*}}{2}$ in equation 38. Combining the above inequality and equation 37, we obtain
$\mathbb{P}\left[W_{t}^{A}<\frac{2}{\alpha}\right]\leq\exp\left(-\frac{(\omega^{A}_{*})^{2}}{2(\ln 3)^{2}}t\right).$ (39)
Similarly, we have
$\mathbb{P}\left[W_{t}^{B}<\frac{2}{\alpha}\right]\leq\exp\left(-\frac{(\omega^{B}_{*})^{2}}{2(\ln 3)^{2}}t\right),$ (40)
where $\omega^{B}_{*}:=\mathbb{E}\left[\ln\left(1-(-g-\epsilon)\theta^{B}_{*}\right)\right]$ for a benchmark $\theta^{B}_{*}$.
Let $H_{t}$ be the event that we stop at time $t$, i.e., $H_{t}:=\{W_{t}^{A}\geq\frac{2}{\alpha}\text{ or }W_{t}^{B}\geq\frac{2}{\alpha}\}$. Then, we deduce that
$\mathbb{P}[t=\infty]$ $=\mathbb{P}\left[\lim_{t\to\infty}\cap_{s\leq t}\neg H_{s}\right]$
$=\lim_{t\to\infty}\mathbb{P}\left[\cap_{s\leq t}\neg H_{s}\right]$
$\leq\lim_{t\to\infty}\mathbb{P}\left[\neg H_{t}\right]$
$=\lim_{t\to\infty}\mathbb{P}\left[W_{t}^{A}<\frac{2}{\alpha}\text{ and }W_{t}^{B}<\frac{2}{\alpha}\right]$
$\leq\lim_{t\to\infty}\mathbb{P}\left[W_{t}^{A}<\frac{2}{\alpha}\right]+\mathbb{P}\left[W_{t}^{B}<\frac{2}{\alpha}\right]$
$\leq\lim_{t\to\infty}\exp\left(-\frac{(\omega^{A}_{*})^{2}}{2(\ln 3)^{2}}t\right)+\exp\left(-\frac{(\omega^{B}_{*})^{2}}{2(\ln 3)^{2}}t\right)$
$=0,$
where the last equality is by equation 39 and equation 40. This shows that the test power is one.
3. Expected Rejection Time $\tau$. Now we upper-bound the expected stopping time as follows. Denote $t_{*}$ the time when equation 39 and equation 40 starts to hold at any $t\geq t_{*}$. We have
$\mathbb{E}[\tau]$ $=\sum_{t=1}^{\infty}\mathbb{P}\left[\tau>t\right]$
$=\sum_{t=1}^{\infty}\mathbb{P}\left[\cap_{s\leq t}\neg H_{s}\right]$
$\leq\sum_{t=1}^{\infty}\mathbb{P}\left[\neg H_{t}\right]$
$\leq\sum_{t=1}^{\infty}\left(\mathbb{P}\left[W_{t}^{A}<\frac{2}{\alpha}\right]+\mathbb{P}\left[W_{t}^{B}<\frac{2}{\alpha}\right]\right)$
$\leq t_{*}+\sum_{t=1}^{\infty}\exp\left(-\frac{(\omega^{A}_{*})^{2}}{2(\ln 3)^{2}}t\right)+\exp\left(-\frac{(\omega^{B}_{*})^{2}}{2(\ln 3)^{2}}t\right)$
$\leq t_{*}+\frac{1}{\exp\left(\frac{(\omega^{A}_{*})^{2}}{2(\ln 3)^{2}}t\right)-1}+\frac{1}{\exp\left(\frac{(\omega^{B}_{*})^{2}}{2(\ln 3)^{2}}t\right)-1}$
$\leq t_{*}+\frac{2(\ln 3)^{2}}{(\omega^{A}_{*})^{2}}+\frac{2(\ln 3)^{2}}{(\omega^{B}_{*})^{2}},$ (41)
where the last inequality uses $\exp(z)\geq 1+z$ for any $z\geq 0$.
To proceed, we recall the notation $\omega^{A}_{*}:=\mathbb{E}\left[\ln(1-(g-\epsilon)\theta_{*})\right]$. We now let the benchmark $\theta_{*}$ to be the one that maximizes the expected wealth in one round, i.e., $\mathbb{E}\left[\ln(1-(g-\epsilon)\theta)\right]$, in the decision space $\mathcal{K}^{\prime}_{*}$.
By the inequality $\ln(1+c)\geq c-c^{2}$ for any $|c|\leq\frac{1}{2}$, we have
$\omega^{A}_{*}=\max_{\theta\in K^{\prime}_{*}}\mathbb{E}\left[\ln\left(1-(g-\epsilon)\theta\right)\right]$ $\geq\max_{\theta\in K^{\prime}_{*}}\mathbb{E}\left[-\theta(g-\epsilon)-\theta^{2}(g-\epsilon)^{2}\right]$
$=\max_{\theta\in K^{\prime}_{*}}-\theta(\mu_{x}-\mu_{y}-\epsilon)-\theta^{2}\left(\mathrm{Var}[g-\epsilon]+(\mu_{x}-\mu_{y}-\epsilon)^{2}\right),$
$\approx\frac{(\Delta-\epsilon)^{2}}{\mathrm{Var}[g-\epsilon]+(\Delta-\epsilon)^{2}}$
$\approx\frac{(\Delta-\epsilon)^{2}}{(d^{\prime}_{*})^{2}+(\Delta-\epsilon)^{2}},$
where we note $\mathbb{E}[g]=\mathbb{E}[\phi(x)-\phi(y)]=\mu_{x}-\mu_{y}$, $\mathrm{Var}[g-\epsilon]=\mathbb{E}[(g-\epsilon)^{2}]-(\mathbb{E}[(g-\epsilon)])^{2}$, and $\mathrm{Var}[g-\epsilon]\leq(d^{\prime}_{*})^{2}$. Similarly, we have $\omega^{B}_{*}\approx\frac{(\Delta-\epsilon)^{2}}{(d^{\prime}_{*})^{2}+(\Delta-\epsilon)^{2}}$.
Combining equation 36 and equation 41, together with the expression of $\omega^{A}_{*}$ and $\omega^{B}_{*}$ above, we have
@@
$\mathbb{E}[\tau]=\mathcal{O}\left(\frac{(d_{*}+\epsilon)^{2}}{(\Delta-\epsilon)^{2}}\ln\left(\frac{d_{*}+\epsilon}{\alpha(\Delta-\epsilon)}\right)+\frac{(d_{*}+\epsilon)^{4}}{(\Delta-\epsilon)^{4}}\right).$ (42)
@@
Appendix G Experimental Results of Detecting 2024 Olympic News or Machine-generated News